You have just made your first $5,000 contribution to your RRSP. Assuming you earn an 11 percent rate of return and make no additional contributions, what will your account be worth when you retire in 45 years? What if you wait ten years before contributing? (Does this suggest an investment strategy?)
Answer #1

Future Value = Present value (1+Interest rate)^Number of Years

Amount after 45 years = 5,000(1.11)^45

= $547,651.21

Amount if contributed after 10 years = 5,000(1.11)^35

= $192,874.26
Answer #2

SOLUTION :


a. Future value starting now =  547,651.21 (ANSWER)


Work :


FV available after 45 years :

= Amount invested ( 1 + r)^n 

= 5000 (1 + 0.11)^45

= 547651.21 ($)  




b.  Future value if waiting 10 years = 192,874.26 (ANSWER)


Work :


If deposit of $5,000 is made after waiting 10 years from now,

FV available after 45 years :

= Amount invested ( 1 + r)^n 

= 5000 (1 + 0.11)^(45 - 10)

= 192874.26 ($) 

answered by: Tulsiram Garg